### v3

# Adaptive Rank Allocation for Federated Parameter- Efficient Fine-Tuning of Language Models

Abstract
Pre-trained Language Models (PLMs) have demonstrated their superiority and versatility in modern Natural Language Processing (NLP), effectively adapting to various downstream tasks through further fine-tuning. Federated Parameter-Efficient Fine-Tuning (FedPEFT) has emerged as a promising solution to address privacy and efficiency challenges in distributed training for PLMs on resource-constrained local devices. However, our measurements reveal two key limitations of FedPEFT: heterogeneous data across devices exacerbates performance degradation of low-rank adaptation, and a fixed parameter configuration results in communication inefficiency. To overcome these limitations, we propose FedARA, a novel Adaptive Rank Allocation framework for federated parameter-efficient fine-tuning of language models. Specifically, FedARA employs truncated Singular Value Decomposition (SVD) adaptation to enhance similar feature representation across clients, significantly mitigating the adverse effects of data heterogeneity. Subsequently, it utilizes dynamic rank allocation to progressively identify critical ranks, effectively improving communication efficiency. Lastly, it leverages rank-based module pruning to automatically remove inactive modules, steadily reducing local computational cost and memory usage in each federated learning round. Extensive experiments show that FedARA consistently outperforms baselines by an average of 6.95% to 8.49% across various datasets and models under heterogeneous data while significantly improving communication efficiency by 2.40$\times$. Moreover, experiments on various edge devices demonstrate substantial decreases in total training time and energy consumption by up to 48.90% and 46.95%, respectively.
Index Terms:
Federated Learning, Communication Efficiency, Data Heterogeneity, Parameter-Efficient Fine-Tuning.
I Introduction
Pre-trained Language Models (PLMs) [1, 2, 3, 4] have demonstrated extraordinary performance and foundational capabilities in modern Natural Language Processing (NLP). Fine-tuning of PLMs further enhances their performance on downstream tasks, such as sequence classification, summarization, and question answering. Pre-training and fine-tuning are evolving into two distinct yet mutually reinforcing stages in NLP. 1) Pre-training: it involves self-supervised training on large-scale corpora from sources like books, news articles, and Wikipedia, aiming to learn general representations of contextual language and create a generalizable pre-trained model. 2) Fine-tuning: it extends the pre-trained model to specific downstream tasks through supervised training, optimizing its performance to meet domain requirements. Overall, these two stages greatly enhance the model’s generality and flexibility across diverse NLP tasks.
Although PLMs have demonstrated superior performance on general-purpose tasks [5, 6], they still require fine-tuning with local data to achieve optimal results in specific domains. However, since this data is typically generated by users, directly collecting it poses a significant risk of privacy leakage. To address this issue, federated learning (FL) is considered an effective approach [7], as it protects data privacy by keeping raw data on local devices, ensuring compliance with regulations such as GDPR [8] and the Data Protection Act [9]. The FL process involves a central server orchestrating the aggregation of local parameter updates from individual clients, with this cycle repeating until training converges.
Figure 1: The complete training process of Federated Parameter-Efficient Fine-Tuning (FedPEFT), where local updates utilize Low-Rank Adaptation (LoRA) while keeping the pre-trained model frozen.
However, the massive scale of PLMs, often with hundreds of millions to billions of parameters [10], incurs substantial computation and communication costs in FL, rendering deployment on resource-constrained devices nearly impossible. To address this issue, a promising solution is to adopt Parameter-Efficient Fine-Tuning (PEFT) methods, which train only a minimal set of parameters using adaptable modules while keeping the pre-trained model frozen [11, 12, 13, 14, 15]. Among these PEFT methods, Low-Rank Adaptation (LoRA) [15] has gained widespread adoption for its effectiveness and efficiency, achieved by introducing two trainable low-rank matrices alongside the pre-trained model for diverse task adaptation. By combining PEFT with FL, distributed data resources on local devices can be fully leveraged for collaborative fine-tuning in a privacy-preserving manner. Consequently, Federated Parameter-Efficient Fine-Tuning (FedPEFT) shows significant advancements and finds wide-ranging applications across various domains [16, 17, 18, 19, 20, 21, 22]. Building on this progress, FedPEFT has increasingly gained recognition as a key method for distributed fine-tuning on local devices [23, 24]. Figure 1 shows the complete FedPEFT training process with LoRA in local clients.
Nevertheless, FedPEFT still faces two major challenges when applied to local devices. 1) FedPEFT exacerbates the impact of non-independent and identically distributed (non-IID) data on model performance. Although performance degradation caused by non-IID data is a well-known challenge in FL, our preliminary measurements show that FedPEFT exacerbates this issue significantly. Specifically, on the 20News [25] dataset using the DistilBERT [2] model, FedPEFT methods exhibit an average accuracy drop of 14.62% under non-IID conditions compared to IID settings—more than double the 6% drop observed with federated full fine-tuning [19]. 2) Fixed rank configurations in FedPEFT incur dual losses in performance and communication cost. Low-rank parameters can be inserted into various positions within the model, such as layers or components like the query, key, and value matrices. However, static allocations often underfit high-sensitivity positions, leading to suboptimal performance—our experiments reveal accuracy gaps of up to 9.76% across different configurations. Meanwhile, redundant parameters assigned to less critical components are repeatedly transmitted during FL rounds without corresponding performance gains, resulting in communication inefficiency.
These dual losses become more severe under non-IID settings, where data heterogeneity intensifies the subspace divergence of federated low-rank adaptation among client-specific models. This divergence leads to greater variability in rank requirements, making fixed rank configurations increasingly misaligned with actual capacity needs. Consequently, both performance degradation and communication overhead are further amplified. Therefore, these two challenges—data heterogeneity and fixed rank allocation—must be jointly addressed to fully unlock the potential of FedPEFT in practical federated learning scenarios.
To address these challenges, we propose Adaptive Rank Allocation for Federated PEFT of language models (FedARA), designed to enhance performance and communication efficiency with minimized resource costs by integrating three novel methods: 1) Truncated SVD adaptation (Section IV-A). We hypothesize that the performance degradation of LoRA arises from its inherent structure and asymmetric initialization. Inspired by Singular Value Decomposition (SVD), we propose a novel truncated SVD adaptation scheme that inserts a diagonal matrix into the low-rank subspace to enable independent feature scaling while preserving symmetric initialization. This mitigates client drift and enhances cross-client feature alignment, significantly alleviating the negative impact of non-IID data. 2) Dynamic rank allocation (Section IV-B). To avoid fixed parameter configurations, we present a dynamic rank allocation method that generates local rank masks through triplets on each client and produces global rank masks via threshold-based arbitration on the server. During federated fine-tuning, less critical ranks are progressively pruned, retaining only significant ones based on global rank masks, effectively improving the communication efficiency of FedPEFT. 3) Rank-based module pruning (Section IV-C). Building on the above two methods, the rank of SVD modules dynamically evolves over FL rounds, we thereby adopt a rank-based module pruning mechanism to automatically remove inactive modules during training. By continuously monitoring rank changes, modules with ranks that fall to zero are safely excluded from the trainable parameters, steadily reducing the computational and storage overhead for client-side training.
Results. We implemented FedARA and evaluated it on three representative edge devices, NVIDIA Jetson AGX Orin, Orin Nano, and Raspberry Pi 5, whose hardware profiles reflect typical resource-constrained settings. We utilize three pre-trained models across five diverse NLP datasets. FedARA outperforms baselines [22, 26, 19, 20] on classification tasks by an average of 6.95% to 8.49% under non-IID data while significantly improving communication efficiency by an average of 2.40$\times$. Remarkably, the total training time on the above devices also decrease by 27.65%, 46.57%, and 48.90%, respectively, and the energy consumption drops by 46.95% on Orin Nano. Ablation experiments reveal that truncated SVD adaptation improves the average accuracy by 7.71%, dynamic rank allocation reduces communication overhead by 70.82% per round while maintaining on-par performance, and rank-based module pruning lowers the average local computation and peak GPU memory usage per round by 10.81% and 31.67%, respectively.
We summarize our key contributions below:
  • We propose FedARA, the first comprehensive approach to address both the adverse effects of non-IID data and fixed parameter configurations in FedPEFT for resource-constrained devices.
  • We create and integrate three innovative methods into FedARA: 1) truncated SVD adaptation strengthens the structure and adjusts initialization to enhance similar feature representation across clients; 2) dynamic rank allocation generates rank masks locally and arbitrates them globally to reduce the communication of less important parameters; 3) rank-based module pruning monitors dynamic rank changes to remove inactive modules.
  • We validate the effectiveness of FedARA through extensive experiments across various datasets and models on advanced edge devices, demonstrating that FedARA consistently improves performance under non-IID data and significantly boosts systems efficiency in both communication and computation.
The rest of the paper is organized as follows: Section II introduces the background of FedPEFT and explains the motivation of our work. Section III presents the proposed FedARA in terms of system architecture and algorithm process. Section IV details the three innovative methods in FedARA. Section V describes the complete experimental setup. Section VI shows and analyzes the experimental results, including ablation studies and extended experiments. Section VII reviews the related work. Finally, Section VIII concludes the paper.
II Background and Motivations
II-A Federated Parameter-Efficient Fine-Tuning (FedPEFT)
(a) FedFFT vs. FedPEFT
(b) The effect of non-IID
(c) The impact of positions
(d) FedPEFT bottlenecks
Figure 2: The preliminary measurement results of FedPEFT in a homogeneous FL setting: (a) Comparison of overall training time on different edge devices (BERT; 20News; Rank=12; Batch size=4). (b) Accuracy of FedPEFT under IID and non-IID scenarios (DistilBERT; Rank=24; Adapter-h (A-h) size=108; Adapter-p (A-p) size=216). (c) Impact of inserted positions on accuracy (DistilBERT; Rank=24). (d) Communication and computation time percentage across models and devices using FedLoRA (News Category; Rank=12; Batch size=4).
FedPEFT orchestrates PEFT strategies within the FL framework to efficiently update local models, drastically reducing communication overhead by only transmitting a small number of PEFT modules. These PEFT modules can be replaced by Adapter [11, 12], Prompt[13, 14], or LoRA [15]. Among them, LoRA-based methods are especially popular because they introduce no additional latency during inference. FedPEFT not only preserves privacy but also effectively aggregates diverse knowledge across clients, facilitating the collaborative development of a high-quality global model. By harnessing the advantages of both PEFT and FL, FedPEFT has become an essential technique for deploying fine-tuning on resource-constrained devices [23, 24, 20, 18, 17, 21].
As illustrated in Figure 1, FedPEFT training process unfolds as follows: Step 1: A cloud server selects a subset of clients to broadcast the initial parameters. Step 2: On these clients, the weights of the pre-trained model are frozen, and trainable PEFT modules are inserted for local training. Step 3: Only these PEFT modules are updated to the cloud server from each client. Step 4: The cloud server aggregates these updates from different clients. This cycle (Steps 1-4) is repeated for multiple rounds in FL until the model converges.
II-B Preliminary Measurements
We conducted a series of preliminary measurements to explain the motivations behind the design of FedARA.
Observation-1: FedPEFT is more effective than FedFFT on advanced edge devices. FedPEFT comprises two main methods: adapter-based [23] and LoRA-based methods [16, 17, 18, 19, 21]. LoRA-based methods have gradually replaced adapter-based methods to avoid additional inference latency. According to established work [23], FedPEFT achieves comparable accuracy to federated full fine-tuning (FedFFT) while significantly reducing local training time by 1.6$\times$, communication cost by 200$\times$, and memory usage by 1.5$\times$ on previous-generation edge devices like NVIDIA Jetson TX2, Nano, and Raspberry Pi 4B. However, its impact on overall training time on advanced devices remains underexplored. To fill this gap, we evaluate FedLoRA on the 20News [25] dataset with the BERT [3] model using latest edge devices. Following FedPEFT’s typical settings [23], we set the communication bandwidth to 1MB/s. As depicted in Figure 2a, the results indicate that FedFFT’s overall training time is 42$\times$, 40$\times$, and 18$\times$ longer than FedLoRA on AGX Orin, Orin Nano, and RPi 5, respectively. FedPEFT demonstrates its effectiveness compared to FedFFT for cutting-edge embedded devices.
Observation-2: The performance of FedPEFT significantly decreases under non-IID data. Figure 2b illustrates the accuracy of different FedPEFT methods using the DistilBERT [2] model for both IID (white bars) and non-IID data (black bars). The non-IID division follows the FedAvg setting [7], with a severe class imbalance in data distribution. The parameter sizes for Adapter-h (A-h) [11] and Adapter-p (A-p) [12] are set to match those of the LoRA method. Our measurements indicate a significant performance drop in all FedPEFT methods under non-IID conditions compared to IID conditions, with average accuracy reductions of 14.62% and 12.81% on the 20News [25] and News Category [27] datasets, respectively. Existing research lacks effective strategies to mitigate the adverse impact of non-IID data on FedPEFT performance.
Observation-3: The contribution of PEFT modules in FedPEFT to performance varies their inserted positions, leading to suboptimal performance and inefficient communication in fixed parameter configurations. In FedPEFT, low-rank modules can be placed at various layers and components, such as Query (Q), Key (K), Value (V), Out (O), F1, and F2. Here, O denotes the linear layer in the multi-head attention block, while F1 and F2 represent the two linear layers of the feed-forward network. As shown in Figure 2c, we allocated a fixed rank budget to specific components or layers to evaluate their impact on accuracy. The left sub-figure shows the allocation of the entire rank budget to a single layer, while the right sub-figure illustrates its distribution across a single component in all layers. The findings reveal that deeper layers near the output often have a more significant effect than shallower ones, with components F1 and F2 typically exerting a greater influence than component O. Moreover, the impact of these inserted positions varies across datasets, suggesting that there is no one-size-fits-all configuration. Fixed rank configurations can result in inefficient communication, as they repeatedly transmit less impactful parameters over hundreds of rounds, yielding minimal performance gains under a limited parameter budget.
Observation-4: Bottlenecks differ across different types of advanced edge devices. In experiments with LoRA-based methods [16, 17] on the News Category [27] dataset, given a communication bandwidth of 1MB/s, we found that the bottlenecks varied depending on the device type, as shown in Figure 2d. On high-end devices, communication is the predominant bottleneck, with communication-to-computation ratios of 4.44 and 5.35 for different models. In contrast, on Raspberry Pi devices, computation becomes the primary bottleneck, with communication-to-computation ratios of 0.67 and 0.72. Importantly, the bottleneck shifts towards communication as model parameters increase, consistent with recent studies [24].
Figure 3: System architecture of FedARA, illustrating the interaction between the server and client, including ClientsSel (Client Selection), CommPru (Communication Pruning), FedArb (Federated Arbitration), MaskGen (Rank Mask Generation), and RankDet (Local Rank Detection).
Insights: These observations underscore key directions for advancing our research. First, given its established advantages, such as improved training efficiency, future research could benefit from moving beyond comparisons with FedFFT and instead focusing on addressing the unique challenges it presents in FL. Second, the performance degradation of different PEFT methods under non-IID remains a critical barrier, necessitating further investigation into effective approaches. Third, fixed parameter configurations lead to communication inefficiency by repeatedly transmitting less impactful parameters over multiple FL rounds, particularly in resource-constrained environments. Finally, optimizing FedPEFT for edge devices requires further refinement, as different hardware presents distinct bottlenecks. To bridge these gaps, enhancing fine-tuning structures and dynamically adjusting parameters could be essential while simultaneously minimizing both communication and computational overhead.
III Overview of FedARA
Due to LoRA’s limitations in handling complex data heterogeneity and the communication inefficiency of fixed parameter configurations, we propose FedARA, an adaptive rank allocation framework for FedPEFT of language models. Our approach introduces a truncated SVD structure to better address data heterogeneity in federated settings. Meanwhile, we incorporate a dynamic rank allocation algorithm to enhance communication efficiency. Additionally, we employ rank-based module pruning to reduce computational time and peak memory usage per round in clients.
The system architecture of the FedARA framework, shown in Figure 3, includes a server and multiple clients. Clients utilize the truncated SVD adaptation to train local models and generate local masks, while the server aggregates the global model and arbitrates global masks. The ranks of the global model are dynamically allocated through the arbitrated global masks. Communication between the server and clients involves transmitting masks and progressively pruned parameters. Local clients detect rank changes in each round to remove inactive SVD modules safely.
Algorithm 1 Federated Adaptive Rank Allocation (FedARA)
1:  Input: all clients $C$, rounds $R$, initial SVD modules $P_{0}$, initial global rank masks $M_{0}$, threshold $T_{h}$, target average rank $T_{r}$, test dataset $D_{t}$, local datasets $D_{l}$, the number of trainable parameters $N_{t}$
2:  Output: optimized SVD modules $P^{*}$, optimized global rank masks $M^{*}$
3:  SERVER:
4:  Initialize $P_{g}\leftarrow P_{0}$, $M_{g}\leftarrow M_{0}$
5:  for $r\leftarrow 1$ to $R$ do
6:     $C_{s}\leftarrow\textbf{ClientsSel}(C)$
7:     // Prune broadcasting parameters by $M_{g}$:
8:     $\hat{P}_{g}\leftarrow\textbf{CommPru}(P_{g},M_{g})$
9:     $\textbf{BroadcastParams}(C_{s},\hat{P}_{g},M_{g})$
10:     // Parallelism among different clients:
11:     $\{\hat{P}_{l+1}\},\{M_{l+1}\}\leftarrow\text{CLIENT\_TRAIN}(r,\hat{P}_{g},M_{g})$
12:     // Aggregation based on FedAvg:
13:     $P_{g+1}\leftarrow\textbf{FedAvg}(\{\hat{P}_{l+1}\},M_{g})$
14:     // Arbitration based on the threshold $T_{h}$:
15:     $M_{g+1}\leftarrow\textbf{FedArb}(\{M_{l+1}\},T_{h},M_{g})$
16:  end for
17:  $P^{*}\leftarrow P_{g},M^{*}\leftarrow M_{g}$
18:  Evaluate final model: $acc\leftarrow\textbf{Eval}(P^{*},D_{t})$
19:  return $acc,P^{*},M^{*}$
20:  CLIENT_TRAIN$(r,\hat{P}_{g},M_{g})$
21:        $P_{l}\leftarrow\hat{P}_{g}$, $M_{l}\leftarrow M_{g}$
22:        $P_{l+1}\leftarrow\textbf{Train}(D_{l},P_{l},N_{t})$
23:        // Generate local rank masks:
24:        $M_{l+1}\leftarrow\textbf{MaskGen}(r,T_{r},P_{l+1})$
25:        // Detect rank changes per round:
26:        $N_{t}\leftarrow\textbf{RankDet}(M_{l+1})$
27:        // Prune uploading parameters:
28:        $\hat{P}_{l+1}\leftarrow\textbf{CommPru}(P_{l+1},M_{l})$
29:        return $\hat{P}_{l+1},M_{l+1}$
Algorithm 1 outlines the process of federated adaptive rank allocation. It starts with parameter initialization and client selection (line 5-7). The server prunes parameters using global rank masks (lines 8-9) and broadcasts them to the clients (line 10). Clients perform parallel local training (lines 11-12), updating parameters based on local datasets (lines 24). Local rank masks are generated from updated parameters (lines 25-26), determining the number of trainable parameters (lines 27-28). During uploading, global rank masks guide parameter pruning (lines 30-31). After receiving client updates, the server aggregates parameters using the FedAvg algorithm [7] (lines 13-14), refines global rank masks via threshold arbitration (lines 15-16), and generates global parameters. Finally, model accuracy is assessed using a test dataset (lines 18-20).
IV Design Details of FedARA
In this section, we present the design of FedARA by elaborating on three interrelated methods: truncated SVD adaptation, dynamic rank allocation, and rank-based module pruning.
IV-A Truncated SVD Adaptation
In previous work, LoRA [15] decomposes the trainable weight matrix into pre-trained weights and incremental weights as:
$W=W_{\text{pre}}+\Delta W=W_{\text{pre}}+\frac{\alpha}{r}BA,$ (1)
where $W_{\text{pre}},\Delta W\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$, $B\in\mathbb{R}^{d_{\text{out}}\times r}$, $A\in\mathbb{R}^{r\times d_{\text{in}}}$, and $\frac{\alpha}{r}$ is a scaling factor with $r\ll\min\{d_{\text{out}},d_{\text{in}}\}$. During training, only the low-rank matrices $A$ and $B$ are updated, while $W_{\text{pre}}$ remains frozen. This decomposition improves parameter efficiency when $r(d_{\text{out}}+d_{\text{in}})<d_{\text{out}}d_{\text{in}}$. Typically, $A$ is initialized with Gaussian noise, and $B$ is initialized to zero, ensuring $\Delta W=0$ at initialization.
However, our pre-experiments and prior studies [19, 20] reveal a significant performance gap in FedLoRA between IID and non-IID settings. We hypothesize that this gap stems from the inherent structure and asymmetric initialization of LoRA. Specifically, LoRA’s low-rank structure offers a high degree of flexibility but inherently couples magnitude and directional updates within the same matrix, amplifying client divergence under data heterogeneity. Additionally, its asymmetric initialization introduces randomness and bias in early training, further intensifying client drift in non-IID settings.
To address this issue, we draw inspiration from the structure of truncated Singular Value Decomposition (SVD), which decomposes a matrix into the eigenvalues and eigenvectors of its low-rank subspace. Accordingly, we propose a truncated SVD adaptation method that mirrors this structure, formulated as
$W=W_{\text{pre}}+\Delta W\;=\;W_{\text{pre}}\;+\;\frac{\alpha}{r}BEA,$ (2)
where $E\in\mathbb{R}^{r\times r}$ is a diagonal matrix, and $B\in\mathbb{R}^{d_{\text{out}}\times r}$, $A\in\mathbb{R}^{r\times d_{\text{in}}}$. Both $A$ and $B$ are initialized with random Gaussian values, while $E$ is initialized to zero, ensuring that $\Delta W=0$ at the beginning of training. In contrast to the standard truncated SVD, we do not impose orthogonality constraints on $B$ and $A$, as such constraints reduce flexibility and incur additional computational overhead in each training round.
Figure 4: Illustration of truncated SVD adaptation, where a diagonal matrix $E$ is inserted between the LoRA matrices.
As illustrated in Figure 4, a diagonal matrix $E$ is inserted between the low-rank subspaces at negligible cost ($r\ll\min(d_{\text{in}},d_{\text{out}})$. The diagonal structure of $E$ restricts updates to independent scaling along each latent dimension, which helps mitigate inter-client variability under non-IID data. At the same time, symmetric initialization ensures balanced gradient flow, offering clients a stable starting point and enhancing global consistency during federated fine-tuning.
We now formally provide theoretical insights into the effectiveness of truncated SVD adaptation under non-IID settings. For each client $k$, let the rank–$r$ update be defined as:
$\Delta W_{k}^{\text{BA}}=\sum_{i=1}^{r}b_{k,i}a_{k,i}^{\top},\quad\Delta W_{k}^{\text{BEA}}=\sum_{i=1}^{r}e_{k,i}b_{k,i}a_{k,i}^{\top},$ (3)
and we quantify client heterogeneity by the drift variance:
$\sigma^{2}=\frac{1}{K}\sum_{k=1}^{K}\!\bigl{\|}\Delta W_{k}-\overline{\Delta W}\bigr{\|}_{F}^{2}\;\approx\;\mathbb{E}\!\left[\!\|\Delta W\|_{F}^{2}\right].$ (4)
Let $T_{i}$ denote the $i$-th rank-1 term. Then, the expression $\|\Delta W\|_{F}^{2}$ can be reformulated as:
$\|\Delta W\|_{F}^{2}=\sum_{i=1}^{r}\sum_{j=1}^{r}\langle T_{i},T_{j}\rangle,$ (5)
where $\langle T_{i},T_{j}\rangle=\operatorname{Tr}(T_{i}^{\top}T_{j})$. Using the identity $\langle uv^{\top},xy^{\top}\rangle=(u^{\top}x)(v^{\top}y)$, the inner product becomes:
$\langle T_{i},T_{j}\rangle=\begin{cases}(b_{i}^{\top}b_{j})(a_{i}^{\top}a_{j}),&\text{(BA)}\\ e_{i}e_{j}\,(b_{i}^{\top}b_{j})(a_{i}^{\top}a_{j}),&\text{(BEA)}\end{cases}\quad,$ (6)
We adopt the standard separable-correlation model:
$\mathbb{E}[b_{i}^{\top}b_{j}]=\begin{cases}\tau_{b},&i=j\\ \rho_{b},&i\neq j\end{cases}\quad,\quad\mathbb{E}[a_{i}^{\top}a_{j}]=\begin{cases}\tau_{a},&i=j\\ \rho_{a},&i\neq j\end{cases}\quad,$ (7)
where $\rho_{a}\rho_{b}\neq 0$ captures non-IID data. For the truncated SVD adaptation, we further assume:
$\mathbb{E}[e_{i}]=0,\quad\mathbb{E}[e_{i}^{2}]=\tau_{e},\quad\bigl{|}\mathbb{E}[e_{i}e_{j}]\bigr{|}\leq\frac{c}{r}\quad(i\neq j).$ (8)
Assumption (8) is justified since each $e_{i}$ is initialized as $e_{i}^{(0)}=0$ and updated by $e_{i}\leftarrow e_{i}-\eta\langle\nabla\ell(x;W),\,b_{i}a_{i}^{\top}\rangle$, where $e_{i}$ only interacts with its own rank-1 component, leading to independent gradient noise across $i\neq j$. Under symmetric mini-batch sampling, the noise is unbiased with $\mathbb{E}[\Delta e_{i}]=0$. Additionally, with Gaussian initialization, the inner product $(b_{i}^{\top}b_{j})(a_{i}^{\top}a_{j})=O(1/\sqrt{d})$, yielding residual correlation of order $O(1/r)$ when $r<\sqrt{d}$.
Substituting equations (6)–(8), we obtain for LoRA:
$\mathbb{E}\!\left[\|\Delta W^{\text{BA}}\|_{F}^{2}\right]=r\tau_{b}\tau_{a}+r(r-1)\rho_{b}\rho_{a}=\Theta(r^{2}),$
while for truncated SVD adaptation, the $r(r-1)$ cross terms are scaled by $\mathbb{E}[e_{i}e_{j}]$ and vanish up to order $O(r)$, yielding:
$\mathbb{E}\!\left[\|\Delta W^{\text{BEA}}\|_{F}^{2}\right]=r\tau_{e}\tau_{b}\tau_{a}+O(r)=\Theta(r).$
Thus, the variance scales as $\sigma^{2}_{\text{BA}}=\Theta(r^{2})$ and $\sigma^{2}_{\text{BEA}}=\Theta(r)$, confirming that truncated SVD adaptation eliminates the quadratic cross-covariance terms and reduces drift growth from $\Theta(r^{2})$ to $\Theta(r)$. This theoretically supports the effectiveness of the newly diagonal matrix $E$ in mitigating non-IID effects.
Figure 5: Magnitude (left) and direction (right) discrepancies between global and local models under different FL structures. FedLoRA and FedSVD correspond to fine-tuned models utilizing LoRA and truncated SVD adaptation, respectively, with a rank of 12.
To further understand the difference in global weight updates between LoRA and truncated SVD adaptation in FL, we introduce magnitude and directional discrepancies between the global model $\boldsymbol{\theta}_{\text{global}}$ and the local models $\boldsymbol{\theta}_{\text{local}}^{(i)}$ across selected clients. The magnitude discrepancy is defined as:
$\text{Mag}=\sum_{i=1}^{K}\left\|\boldsymbol{\theta}_{\text{global}}-\boldsymbol{\theta}_{\text{local}}^{(i)}\right\|_{F},$ (9)
where $K$ denotes the number of selected clients per round, and $\|\cdot\|_{F}$ represents the Frobenius norm. This metric captures the magnitude difference between the global and local models. The directional discrepancy is given by:
$\text{Dir}=\frac{1}{{K}}\sum_{i=1}^{{K}}\frac{\langle\boldsymbol{\theta}_{\text{global}},\boldsymbol{\theta}_{\text{local}}^{(i)}\rangle}{\|\boldsymbol{\theta}_{\text{global}}\|_{F}\,\|\boldsymbol{\theta}_{\text{local}}^{(i)}\|_{F}}.$ (10)
This metric measures the averaged cosine similarity between the global and local models, with values closer to 1 indicating higher directional alignment.
We fine-tune the DistilBERT [2] model on the 20News [25] dataset in an FL setting with 100 rounds, selecting 10 clients per round. As a representative, we extract the delta weights of layer.3.ffn.lin1, which reconstructs the low-rank module into an incremental matrix matching the dimensions of the pre-trained model. As shown in Figure 5, FedSVD shows greater consistency in global model updates compared to FedLoRA under non-IID settings. FedLoRA exhibits significant sensitivity to heterogeneity, manifesting in pronounced magnitude and direction discrepancies, while FedSVD demonstrates much smaller fluctuations. This highlights that the truncated SVD structure is more effective than LoRA in representing similar features across diverse clients.
IV-B Dynamic Rank Allocation
The fixed rank configuration assigns uniform ranks to modules across different layers and components, resulting in communication inefficiency due to its inability to adapt to the varying contributions of layers and components to model performance. This inefficiency manifests in two aspects: the fixed rank allocation yields suboptimal performance under similar communication overhead and requires higher communication costs to achieve comparable performance with the dynamic rank allocation.
Inspired by AdaLoRA [28], which introduces a framework for centralized rank assignment, we propose a dynamic rank allocation method in FL that efficiently utilizes limited parameters to prioritize the most critical weights, thereby enhancing communication efficiency. This method progressively identifies the most significant weights while pruning less important parameters across clients, involving three key steps: Generation of Local Rank Masks, Arbitration of Global Rank Masks, and Communication with Arbitrated Masks.
Figure 6: Construction of triplets in SVD modules and the structure of local rank masks.
IV-B1 Generation of Local Rank Masks
During federated fine-tuning, local rank mask generation is directly influenced by the rank budget, defined as the total sum of ranks across all layers and components. This budget is progressively reduced, eventually approaching the pre-set average target rank. The relationship between the budget $b^{(t)}$ and the FL round $t$ is given by:
$b^{(t)}=\begin{cases}b^{(0)},&0\leq t<t_{w},\\ b^{(T)}+b^{(c)}\left(1-\frac{t-t_{w}-t_{f}}{T-t_{w}-t_{f}}\right)^{3},&t_{w}\leq t<T-t_{f},\\ b^{(T)},&\text{otherwise}.\end{cases}$ (11)
where $b^{(t)}=b^{(0)}-b^{(R)}$, $t_{w}$ represents the warm-up rounds, $t_{f}$ indicates the final stabilized rounds, $T$ denotes the total rounds.
Local clients identify relatively important rank candidates based on the reduced rank budget. The importance of each rank is evaluated using triplets, as shown in Figure 6, which consist of the diagonal value in $E$, the left vector in $B$, and its corresponding right vector transpose in $A$. This can be represented as:
$I_{n,i}=I(E_{n,i})+\frac{1}{d_{1}}\sum_{j=1}^{d_{1}}I(B_{n,ji})+\frac{1}{d_{2}}\sum_{j=1}^{d_{2}}I(A_{n,ij}),$ (12)
where $I(E_{n,i})$, $I(B_{n,ji})$, and $I(A_{n,ij})$ denote the importance scores of the corresponding parts. As shown in Table I, we compare four importance score evaluation strategies: Mag denotes the weight magnitude, $I(w_{i,j})=|w_{i,j}|$; Grad represents the gradient, $|\text{grad}(w_{i,j})|$; Mixed combines both weight and gradient information via their product, $|\text{grad}(w_{i,j})\cdot w_{i,j}|$; and Sensitivity corresponds to the sensitivity-based method proposed in AdaLoRA [28]. Among the evaluated strategies, the magnitude-based method consistently achieves the best performance across different datasets and models, outperforming both the mixed strategy and gradient-based approaches. While the sensitivity-based method also improves accuracy, it introduces approximately 30% additional computational overhead. To avoid the high cost of sensitivity and gradient-based estimation, we adopt the magnitude-based score as a simple yet effective alternative.
TABLE I: Performance and local computation overhead under different importance score evaluation methods (non-IID alpha=0.1). Results are reported in percentage (%).
Methods DistilBERT BERT Comp.
20News Semeval 20News Semeval
Mag 63.89 80.24 64.32 76.89 1$\times$
Grad 63.31 79.59 63.68 76.05 0.97$\times$
Mixed 63.79 79.87 63.94 76.42 1.11$\times$
Sensitivity 63.89 80.06 64.21 76.70 1.30$\times$
[…]
TABLE II: Performance and communication overhead under different arbitration strategies (non-IID alpha=0.1). Results are reported in percentage(%).
Methods 20News Comm.
DistilBERT BERT
FedARA 62.94 63.52 8.80 GB
FedARA-global 62.78 63.36 8.82 GB
IV-B2 Arbitration of Global Rank Masks
After collecting all candidates of local rank masks from all selected clients, the server determines global rank masks by arbitrating the fraction of clients reporting True at each position against a preset threshold $T_{h}$. This arbitration process is given by:
$M_{\mathrm{global}}^{(i)}=\begin{cases}True,&\text{if }\displaystyle\frac{1}{\lvert{K}\rvert}\sum_{k\in{K}}M_{k}^{(i)}>T_{h},\\ False,&\text{otherwise}.\end{cases}$ (13)
here, ${K}$ denotes the set of participating clients. The arbitrated global rank masks are broadcast to all clients and used to optimize communication parameters between the server and clients. As shown in Table II, we compare two arbitration strategies: FedARA collects local rank masks from clients and performs arbitration on the server; and FedARA-global generates global rank masks directly from the aggregated model. Although the communication overhead is similar for both strategies, FedARA achieves better performance on the 20News dataset [25], highlighting the effectiveness of leveraging local rank information for centralized arbitration.
IV-B3 Communication with Arbitrated Masks
The global and local rank masks are transmitted alongside the parameters, yet their communication overhead remains negligible due to their boolean nature and significantly smaller size relative to the full parameters. During server-client communication, if a global rank mask in the truncated SVD modules is marked as False, the corresponding parameters are pruned. The complete weight matrix is reconstructed on the client or server using global rank masks. This approach substantially reduces communication overhead in each round.
FedARA progressively achieves adaptive rank allocation during federated fine-tuning rounds by following the aforementioned steps, ultimately preserving only the most important ranks while removing less significant ones. This process significantly improves communication efficiency.
IV-C Rank-based Module Pruning
Although the dynamic parameter allocation method with truncated SVD modules reduces communication overhead in FL by limiting the upload and download of less critical ranks, it does not decrease the number of trainable parameters during local training. This limitation arises because frameworks such as PyTorch and Transformers manage trainable parameters at the matrix level, preventing selective updates at the rank level. Fine-grained sparse gradient updates require additional computation overhead. To address this, we introduce a rank-based module pruning mechanism that safely removes inactive SVD modules in a structured manner.
Specifically, each SVD module on the local client has a rank parameter that changes over time according to local training and the ranking results of triplets. Dynamic rank allocation ensures that the ranks of SVD modules either remain constant or gradually decrease as FL training progresses. Critical ranks are preserved, while less significant ones are progressively pruned. Once a rank is reduced to a predefined threshold (set to zero in this work), it is deemed non-essential and can be safely removed.
Thus, FedARA continuously monitors the rank changes of each inserted SVD module in every FL round. When a module’s rank falls below the threshold, all three matrices in the SVD module become non-trainable. This process does not affect model performance, as these inactive ranks are excluded from the communication and aggregation. By eliminating unnecessary training of specific SVD modules, this method reduces computational overhead and memory consumption, making it particularly advantageous for deployment on resource-constrained devices.
V Experimental Setups
FedARA extends AdaLoRA [28], originally designed for dynamic parameter configuration in centralized training, to the federated learning setting. Built on Hugging Face’s Transformers library, it supports a broad range of pre-trained NLP models. FedARA adopts a random client selection mechanism, which can be substituted with more advanced strategies [29, 30]. Communication between clients and the server is synchronized via weighted averaging [7], maintaining compatibility with diverse optimization and aggregation methods [31, 32, 33, 34, 35, 36]. Moreover, FedARA incorporates model quantization techniques such as QLoRA [37] and QA-LoRA [38] to reduce computational and memory overhead, potentially lowering communication costs.
Metrics. The primary evaluation metrics are final accuracy for classification tasks, ROUGE scores [39] for summarization tasks, and total communication overhead after a fixed number of FL rounds, ensuring fair performance comparison. Secondary metrics include per-round communication cost, adaptive rank allocation, total training time, memory footprint, and energy consumption on edge devices. These metrics offer a comprehensive evaluation of system efficiency and effectiveness.
Hardware. Following prior FL research [23], all experiments are conducted via emulation on a laptop equipped with an RTX 4070 GPU. Local training time is measured on three edge devices—AGX Orin, Orin Nano, and Raspberry Pi 5—chosen for their representativeness of common resource-constrained hardware. These results are integrated into our emulation framework to estimate total training time in a homogeneous FL setting. In line with previous studies [23], we assume a communication bandwidth of 1MB/s between the server and clients, reflecting realistic conditions across IoT, home WiFi, and cellular networks.
TABLE III: Different dataset settings and sizes in sequence classification and sumarization tasks of NLP.
Tasks Datasets Client Num. Labels Samples
Classification 20News [25] 100 20 18.8k
Semeval [40] 100 19 10.7k
AG News [41] 1000 4 127.6k
News Category [27] 1000 15 148.1k
Summarization CNN/DailyMail [42] 1000 - 287k
Datasets and Division. FedARA is evaluated on sequence classification and summarization tasks using five diverse NLP datasets, as summarized in Table III. Classification tasks include: (1) 20News [25], comprising documents from 20 different newsgroups; (2) Semeval [40], covering various NLP and semantic evaluation tasks; (3) AG News [41], containing news articles from four categories; (4) News Category [27], a dataset of diverse news headlines, where we focus on the top 15 most frequent labels, covering approximately 70% of the data. Summarization task: (5) CNN/DailyMail [42], a widely used dataset for abstractive summarization. Each classification dataset is split into training, validation, and test sets in an 8:1:1 ratio. We fine-tune a DistilBERT model on the News Category dataset (top-15 labels) and use the resulting model to generate pseudo-labels for the summarization dataset. To simulate non-IID conditions in a federated setup, we adopt Dirichlet-based partitioning with $\alpha\in\{1,0.1,0.01\}$. Following FedAvg [7], we also consider a pathological non-IID setting where each client holds samples from only 1–2 labels. For IID settings, we set $\alpha=1000$ to approximate uniform label distribution across clients.
Model. We adopt three widely used transformer-based models for NLP tasks. For classification, we use BERT-base [3], which consists of 12 transformer layers, and its compressed variant DistilBERT-base [2], which retains most of BERT’s performance with only 6 layers via knowledge distillation. For summarization, we employ BART-base [4], an encoder–decoder model designed for sequence-to-sequence generation. All selected models are based on the transformer architecture, which remains foundational for even the latest NLP models [10]. Our work focuses on federated adaptive rank allocation for transformer models. While this study evaluates BERT, DistilBERT, and BART across classification and summarization tasks to ensure generality and feasibility on resource-constrained devices, our method is readily applicable to newer architectures and broader NLP tasks.
TABLE IV: Comparison of FedPEFT methods across different models and datasets. In A (B), A represents the accuracy under non-IID data (Pathological distribution), while B denotes the relative accuracy degradation compared to IID data.
Models Methods PEFT Comm Overhead (GB) 20News (%) Semeval (%) AG News (%) News Category (%)
DistilBERT FedLoRA (r=6) 11.12 50.81 (↓14.19) 34.76 (↓45.71) 86.74 (↓1.64) 56.34 (↓14.67)
FedAdapter-h (a=27) 11.34 51.97 (↓13.06) 19.09 (↓62.57) 85.93 (↓2.60) 59.29 (↓12.03)
FedAdapter-p (a=54) 11.24 52.19 (↓12.76) 20.04 (↓60.93) 86.33 (↓2.56) 59.47 (↓11.99)
SLoRA (r=6) 11.12 48.83 (↓15.33) 31.74 (↓47.01) 85.68 (↓2.17) 55.57 (↓14.91)
FeDeRA (r=6) 11.12 52.28 (↓11.43) 28.33 (↓50.95) 86.22 (↓1.94) 56.45 (↓14.64)
FFA-LoRA-dr (r=12) 11.12 53.96 (↓10.99) 31.91 (↓48.46) 87.34 (↓1.55) 61.58 (↓10.14)
FFA-LoRA (r=6) 5.56 53.52 (↓11.24) 29.80 (↓50.29) 87.00 (↓1.70) 61.27 (↓10.36)
FedARA (init r=12) 9.29 55.05 (↓10.03) 56.68 (↓24.29) 87.40 (↓1.37) 63.47 (↓8.10)
FedARA (init r=6) 4.64 54.86 (↓10.22) 55.77 (↓25.20) 87.27 (↓1.50) 63.56 (↓8.01)
BERT FedLoRA (r=6) 22.24 46.27 (↓18.03) 34.91 (↓45.35) 86.47 (↓2.44) 55.70 (↓16.09)
FedAdapter-h (a=27) 22.68 49.55 (↓15.31) 39.20 (↓41.15) 86.40 (↓2.52) 56.34 (↓15.50)
FedAdapter-p (a=54) 22.46 49.36 (↓16.04) 31.07 (↓49.14) 85.66 (↓3.41) 57.68 (↓15.14)
SLoRA (r=6) 22.24 45.60 (↓16.73) 28.42 (↓47.93) 84.36 (↓3.37) 54.20 (↓15.76)
FeDeRA (r=6) 22.24 50.58 (↓13.95) 31.69 (↓49.02) 86.97 (↓2.42) 54.82 (↓17.72)
FFA-LoRA-dr (r=12) 22.24 54.77 (↓9.15) 46.51 (↓34.16) 87.39 (↓1.68) 65.27 (↓6.91)
FFA-LoRA (r=6) 11.12 53.10 (↓11.33) 38.25 (↓42.23) 86.67 (↓1.45) 62.60 (↓9.81)
FedARA (init r=12) 18.86 55.61 (↓9.45) 48.67 (↓31.03) 87.40 (↓1.27) 65.48 (↓6.41)
FedARA (init r=6) 9.46 55.08 (↓9.98) 48.35 (↓31.35) 87.32 (↓1.35) 65.71 (↓6.18)
Baselines. We compare FedARA against several recent FedPEFT methods:
  • FedLoRA: a widely adopted baseline with extensive variants and empirical studies [16, 17, 21, 18].
  • FedAdapter-h: incorporates Adapter-h [11] modules into the attention and feed-forward layers of transformers.
  • FedAdapter-p: inserts adapter modules only in the feed-forward layers [12]. We exclude AdaFL [23] as it is a IIIparameter-reduced version of FedAdapter-p without targeted non-IID optimization.
  • SLoRA: employs a two-stage training scheme [19], where the first stage applies sparse fine-tuning (SFT) to improve initialization for LoRA in the second stage. We allocate 10% of the total FL rounds to the SFT stage.
  • FeDeRA: initializes LoRA using singular value decomposition of pre-trained weights [20].
  • FFA-LoRA: updates only the $B$ matrix in LoRA while freezing $A$ [26, 22, 43]. To mitigate reduced expressiveness, we also evaluate FFA-LoRA-dr, which employs orthogonal initialization for $A$ and doubles the rank.
Hyperparameters. Learning rates are selected via grid search in the range of 1e-5 to 5e-3, depending on the dataset and model. All methods use a batch size of 4, one epoch per FL round, a maximum sequence length of 64, and are fine-tuned for 100 rounds with 10 clients per round (10% or 1% selection). For summarization, we set the source length to 512, the target length to 64, and the beam size to 1 to fit edge device constraints. Adaptive rank allocation starts after 5 warm-up rounds and decays until round 50, targeting one-quarter of the original rank with a global mask threshold of 0.5. Learning rates decay linearly across rounds, using the Adam optimizer. The LoRA scaling factor $\alpha$ is fixed at 16, following prior work [15, 28].
VI Evaluation
VI-A Performance under non-IID data
We conduct experiments on classification tasks under pathological non-IID and IID settings, using two different models across four datasets. The results are summarized in Table IV, where the best performance is highlighted in bold and the second-best is underlined. Each result is averaged over five runs with different random seeds.
FedARA effectively mitigates performance degradation under non-IID settings. FedARA consistently outperforms existing methods across four datasets and two models. Using DistilBERT, FedARA achieves an average accuracy improvement of 8.49% over FedLoRA and 11.58% over FedAdapter-h, while also surpassing FFA-LoRA-dr and FFA-LoRA by 6.95% and 7.75%, respectively. The largest gain 37.59% is observed on the Semeval dataset, which has more labels and fewer samples, whereas the smallest gain 1.47% occurs on AG News, with fewer labels and more samples. Moreover, FedARA significantly outperforms SLoRA and FeDeRA, which rely on modified LoRA initialization strategies.
We further conduct experiments under varying degrees of non-IID data distributions by adjusting the Dirichlet parameter $\alpha$, where smaller $\alpha$ indicates higher non-IIDness. The results are illustrated in Figure 7. As $\alpha$ decreases, the performance of FedPEFT shows a clear downward trend, which is consistent with the observations reported in [19, 20]. Moreover, FedARA consistently outperforms other FedPEFT methods across different $\alpha$ values, with the performance gap becoming particularly significant under highly skewed non-IID settings.
Within our experimental framework, SLoRA and FeDeRA underperform FedARA on certain non-IID tasks and datasets. SLoRA requires careful hyperparameter tuning during its initial sparse fine-tuning stage, especially under varying non-IID data distributions and tasks. Its advantage tends to diminish as LoRA’s rapid adaptation under the Adam optimizer. The effectiveness of FeDeRA relies on alignment between the pre-trained model and the downstream task, which becomes increasingly difficult to maintain under non-IID conditions, potentially leading to performance degradation when such alignment is weak. In contrast, FedARA leverages a scaling vector and symmetric initialization to reduce client drift and promote consistent feature representations, thereby mitigating performance degradation.
Figure 7: Accuracy comparison under different data distributions, with parameter settings identical to those in Table IV.
FedARA significantly improves communication efficiency. FedARA reduces the communication overhead while maintaining superior performance. As shown in Figure 8, FedARA begins with a similar initial communication cost (r=12) as baseline methods on the BERT model and AG News dataset, but quickly drops it to 131.42 MB—71.17% lower than its initial warm-up cost. SLoRA and FeDeRA incur the same communication cost as FedLoRA. Table IV further shows that even when FedARA’s initial rank is twice that of FedLoRA, its total communication overhead remains nearly 20% lower than that of FedLoRA and FFA-LoRA-dr. When initial ranks are equal, FedARA reduces communication overhead by approximately 2.40$\times$. FFA-LoRA also incurs about 20% more overhead than FedARA. Overall, FedARA achieves the lowest communication cost across settings while consistently outperforming all baselines.
Figure 8: Comparison of communication overhead per round across baselines.
Figure 9: Results of final adaptive rank allocation using the DistilBERT model.
Figure 9 shows the final rank distribution of FedARA on the DistilBERT model and 20News dataset. Starting from a uniform rank of 12, FedARA adaptively reduces the average rank to 3, effectively balancing performance and communication cost. Darker blocks in Figure 8 represent higher ranks, mainly in deeper layers and key components like query, key, f1, and f2, highlighting their role in preserving important model information—consistent with our earlier findings (Figure 2c).
To evaluate the generalization and scalability of FedARA, we conduct summarization experiments using the BART model on the CNN/DailyMail dataset. FedARA demonstrates strong performance while achieving the best communication efficiency among all compared methods. As shown in Table V, FedARA improves ROUGE scores over FFA-LoRA, FFA-LoRA-dr, and FedLoRA. FedARA delivers performance comparable to SLoRA and FeDeRA while reducing communication overhead by about 60%. These results confirm its effectiveness in generative NLP tasks with encoder-decoder architectures, in addition to classification.
TABLE V: Results on text generation tasks fine-tuned by the BART [4] model for different baselines. non-IID alpha is set to 0.1.
Datasets Methods Comm. (GB) ROUGE 1/2/L (%)
CNN /DailyMail FedLoRA (r=12) 27.19 40.79/18.33/28.70
SLoRA (r=12) 27.19 40.90/18.51/28.98
FeDeRA (r=12) 27.19 40.90/18.49/29.00
FFA-LoRA-dr (r=24) 27.19 40.71/18.28/28.66
FFA-LoRA (r=12) 13.60 40.58/18.14/28.55
FedARA (init r=12) 9.77 40.93/18.49/29.00
Figure 10: Total training time across three different edge devices, including RPi 5, Orin Nano, and AGX Orin, each with varying hardware constraints. The local training time is measured on the DistilBERT model to assess training time differences across these devices.
VI-B Training Time on Advanced Edge Devices
We evaluate the training efficiency of FedARA and baselines on various edge devices. On Raspberry Pi 5 (RPi 5), local training time per batch (batch size = 4) is 1.00 s for DistilBERT and 2.01 s for BERT. AGX Orin and Orin Nano achieve significantly faster training under DistilBERT, with 6.67$\times$ and 5.56$\times$ speedups over RPi 5, respectively; for BERT, the speedups increase to 8.74$\times$ and 6.70$\times$.
Figure 11: Effectiveness of SVD modules in mitigating non-IID problems and the performance of different initial rank configurations for FedARA.
Figure 10 compares total training time for the DistilBERT model across devices. On RPi 5, FedARA reduces training time to 68 minutes, representing 27.65% and 24.38% reductions over FedLoRA and FFA-LoRA-dr, respectively. On Orin Nano and AGX Orin, FedARA further lowers training time to 25 and 23 minutes, achieving up to 48.90% reduction over FedLoRA and over 12% compared to FFA-LoRA. These results suggest that limited compute on edge devices may diminish the benefits of reduced communication overhead. This observation aligns with Figure 2d, where high-end devices are communication-bound, while low-end devices remain compute-bound. We exclude comparisons with SLoRA and FeDeRA due to their noticeably higher training and computational costs compared to FedLoRA. Specifically, SLoRA incurs overhead comparable to full fine-tuning when using optimizer masking approach [18], while FeDeRA introduces additional model-size-dependent costs through SVD.
VI-C Ablation Experiments
To validate the effectiveness of our proposed methods, we conduct ablation studies comparing FedLoRA, FedARA, and FedSVD, where FedSVD replaces LoRA modules with SVD modules but lacks the adaptive rank strategy. All methods are tested with a uniform initial rank of 4 or 8.
Ablation study of truncated SVD adaptation. We first assess the impact of SVD modules on performance. As depicted by the first two bars of the sub-figure in Figure 11, FedSVD improves accuracy by an average of 7.71% across two different models compared to FedLoRA, highlighting the importance of the SVD structure in mitigating performance degradation under non-IID settings.
Ablation study of dynamic rank allocation. To examine the role of dynamic rank allocation, we evaluate FedARA-r8 and FedARA-r4, which have initial ranks of 8 and 4, respectively. As shown in Figure 11, FedARA-r8 improves average performance by 0.45% and reduces total communication overhead by nearly 20% compared to FedSVD. Meanwhile, FedARA-r4 achieves comparable or superior performance while reducing communication overhead by approximately 2.20$\times$. These results confirm that dynamic rank allocation maximizes communication efficiency.
Additionally, we assess the per-round communication overhead across FedLoRA, FedSVD, and FedARA. As illustrated in Figure 12, all methods begin with a similar overhead of around 75.98 MB. While FedLoRA and FedSVD maintain constant communication costs, FedARA’s dynamic rank allocation progressively reduces costs, stabilizing at 22.17 MB after 50 rounds—a 70.81% and 70.82% reduction compared to FedLoRA and FedSVD. This suggests that FedARA’s fine-tuned model requires fewer trainable parameters, offering potential benefits for minimizing storage costs of PEFT modules for various downstream tasks.
Ablation study of rank-based module pruning. Rank-based module pruning does not impact performance, as it preserves the same parameters for communication and global aggregation. We evaluate its effect on local training time and GPU usage on Semeval and News Category datasets. As illustrated in Figure 13 and 14, while FedARA initially exhibits a slight increase in local training time, it significantly reduces over time. Compared to FedLoRA, FedARA lowers average local training time by 10.81% and reduces GPU memory consumption by 31.67% per round, demonstrating its efficiency in minimizing computational and storage overhead.
Figure 12: Effectiveness of dynamic rank allocation in improving communication efficiency per round. (Model: DistilBERT; Dataset: 20News).
VI-D Extended Experiments
To assess the impact of the target average rank ($T_{r}$) and arbitration threshold ($T_{h}$), we conduct extended experiments with an initial rank of 12 and a default threshold of 0.5. As shown in the upper part of Figure 15, lower target ranks reduce communication overhead but compromise validation loss, while higher ranks enhance performance at the expense of increased overhead. To balance performance and communication efficiency, we set the predefined target average rank ($T_{r}$) to one-fourth of the initial rank in this study.
Figure 13: Effectiveness of rank-based module pruning in reducing local training time per round.
Figure 14: Effectiveness of rank-based module pruning in lowering memory usage per round.
Figure 15: Impact of target rank and arbitration threshold on validation loss and communication overhead. The upper subfigure shows results on 20News, and the lower on AG News using the DistilBERT model.
Additionally, the lower part of Figure 15 shows that increasing the arbitration threshold ($T_{h}$) slightly increases validation loss while marginally reducing communication overhead. This is because a higher $T_{h}$ enforces stricter arbitration, leading to more ranks being deemed unimportant. Notably, FedARA is relatively insensitive to $T_{h}$, indicating that local rank masks tend to converge across clients. Based on these observations, we adopt 0.5 as the default value for $T_{h}$ in this study.
VI-E Local Resource Utilization
Memory Footprint. Although SVD adaptation introduces an additional diagonal matrix, the memory overhead remains negligible due to the small rank $r$ relative to the hidden layer size. As shown in Figure 16, integrating SVD modules and dynamic rank allocation in FedARA leads to only a slight increase in peak memory usage compared to FedLoRA—2.42% for DistilBERT and 4.40% for BERT on Orin Nano. Moreover, FedARA progressively reduces memory usage per round during federated fine-tuning, confirming its suitability for memory-constrained devices.
Energy Consumption. Figure 17 shows the energy consumption on Orin Nano (15W) over 100 rounds with 10 clients per round. FedARA achieves the lowest energy usage—approximately $2.3\times 10^{5}$ J for DistilBERT and $4.4\times 10^{5}$ J for BERT—representing a 46.95% reduction compared to FedLoRA and around 9% compared to FFA-LoRA. These results demonstrate FedARA’s ability to reduce energy consumption, extending device runtime and lowering deployment costs for sustainable use in resource-constrained environments.
VII Related Work
FedARA builds on recent advances in Parameter-Efficient Fine-Tuning (PEFT) and its extension to the federated PEFT (FedPEFT). This section reviews key developments in both areas and examines related work addressing two main challenges in FedPEFT: mitigating non-IID effects and supporting dynamic parameter configurations.
PEFT. PEFT enables efficient adaptation to downstream tasks by fine-tuning only a small subset of parameters, significantly reducing resource consumption compared to full fine-tuning. Based on insertion positions and integration mechanisms, PEFT methods can be categorized into: 1) adapter-based methods [11, 12], which insert small modules after transformer linear layers; 2) prompt-based methods [13, 14], which embed trainable parameters into input prompts; and 3) LoRA-based methods [15], which inject low-rank matrices in parallel with transformer linear layers. FedARA adopts LoRA-based modules for local training, allowing seamless integration into PLMs without additional inference latency.
FedPEFT. FedPEFT extends PEFT to federated learning, leveraging its efficiency to reduce the substantial communication overhead of applying PEFT to PLMs. FedIT [16] facilitates privacy-preserving acquisition of high-quality instruction data for generative tasks. FLoRA [17] proposes a noise-free aggregation strategy tailored to heterogeneous FL, while FlexLoRA [18] presents an efficient framework for multi-task FL under resource heterogeneity. Other efforts focus on system-level optimization of FedPEFT [23, 24] and open-source frameworks for LLMs and heterogeneous edge environments [44, 45]. These works offer valuable context for real-world federated LLM deployment and complement our focus on efficient and adaptive federated fine-tuning under non-IID conditions.
Non-IID Data. Non-IID data in FL hinders model aggregation and global generalization. Traditional methods—such as client selection, optimization, and aggregation—have been effective in mitigating these effects [29, 30, 31, 32, 33, 34, 35, 36]. Recent work extends this line to pre-trained models within the FedPEFT framework. SLoRA [19] adopts a two-stage strategy: sparse fine-tuning for initialization, followed by LoRA adaptation. FeDeRA [20] enhances LoRA initialization via SVD. Beyond initialization, recent studies explore asymmetric training by updating only matrix $B$ while freezing matrix $A$ to reduce non-IID impact [43, 26, 22]. FedARA further addresses non-IID challenges by modifying both the structure and initialization via truncated SVD adaptation, effectively slowing client drift.
Figure 16: Comparison of memory footprint on the Orin Nano device.
Figure 17: Comparison of energy consumption on the Orin Nano device.
Dynamic Parameter Configurations. Fixed rank configurations lead to dual losses of suboptimal performance and communication inefficiency. In centralized training, AdaLoRA [28] introduces a sensitivity-based dynamic rank allocation method. In the context of FedPEFT, AdaFL [23] expands parameters via sideline trials on other clients, but introduces inference latency and additional communication costs. SA-FedLoRA [21] employs a two-stage simulated annealing approach: full fine-tuning with regularization to control model drift, followed by adaptive LoRA rank adjustment. However, the high cost of full fine-tuning limits its practicality on resource-constrained devices. FedARA replaces full fine-tuning with truncated SVD adaptation for resource efficiency, and leverages rank masks to reduce communication without extra clients, enabling a more scalable solution.
VIII Conclusion
FedARA is a novel framework for Federated Parameter-Efficient Fine-Tuning (FedPEFT) of language models. It improves performance under non-IID data and improves communication efficiency by integrating adaptive rank allocation with truncated SVD adaptation, while minimizing per-round local resource usage. Extensive experiments validate its superiority over existing baselines. In future work, we aim to extend FedARA to larger models and broader tasks, and explore integration with quantization to further improve adaptability on resource-constrained devices.